OpenStudy (anonymous):

A water balloon has a diameter of 6. How many cubic inches of water does it contain, when it is 3/4 full?

OpenStudy (apoorvk):

Diameter is 6, so what's the radius?

OpenStudy (apoorvk):

Nah, that's not radius -.- Radius is half of the diameter. What say now?

OpenStudy (anonymous):

it would b 3

OpenStudy (anonymous):

your choices r 48pie, 36pie, 6pie, 27pie. or 216pie.

OpenStudy (anonymous):

Avoid wasting your brain tissue by simply pouring the water into a measuring jar and then noting the reading.

OpenStudy (apoorvk):

Yo! so radius = r = 3 Now. Am assuming that the balloon is sphere shaped. Now, Volume of a sphere = \(\large \frac43 \pi r^3\), where 'r' is the radius. So what would be the complete volume of your sphere?

OpenStudy (anonymous):

IDK

OpenStudy (apoorvk):

just plug in r =3 in the equation man. You can do this much!

OpenStudy (apoorvk):

actually you would also multiply the volume by 3/4 later on, since the balloon is only 3/4th filled with water. So, Volume of water in the balloon = \(\large \frac34\times \frac 43 \pi r^3 = \large \frac{\cancel 3}{\cancel4}\times \frac {\cancel4}{\cancel3} \pi r^3 = \pi r^3 \) Now this is dead simple man, jut replace 'r' with '3', and then calculate a step!